OpenStudy (anonymous):

A Growing City: The population of a city grows at a rate of 5 percent per year. The population in 1990 was 400,000. In what year will the population reach 1 million?

OpenStudy (anonymous):

F=P(1+r)^t F: final population P: initial population r: rate t: time

OpenStudy (anonymous):

wait i need step cuz so far i did the exponential growth

OpenStudy (anonymous):

Ignore nincompoop. This is not linear growth.

OpenStudy (anonymous):

which is 791972.6398

OpenStudy (kropot72):

\[10^{6}=4\times 10^{5}(1+0.05)^{x}\] where x is the number of years from 1990. Solve the equation for x.

OpenStudy (anonymous):

a. What is the predicted population in the year 2004? 791972.6398

OpenStudy (anonymous):

sorry the preceding was included as well

OpenStudy (anonymous):

i have no clue how to get the 1 million

OpenStudy (anonymous):

im wondering if i use logs in this

OpenStudy (kropot72):

@yomamabf Have you checked my equation. BTY one million is\[10^{6}\]

OpenStudy (kropot72):

@yomamabf Yes, I solved it using logs.

OpenStudy (anonymous):

yes but i dont think u applied the recent post ... oh u did ok lemme try and tell me if im correct

OpenStudy (kropot72):

Sure. I'll wait.

OpenStudy (anonymous):

actually im really bad at logs and cant figure this out can u explain the steps to me

OpenStudy (anonymous):

log=log400,000(1+.05)^600,000

OpenStudy (anonymous):

? im not good at logs

OpenStudy (kropot72):

\[10^{6}=4\times 10^{5}(1.05)^{x}\] \[(1.05)^{x}=\frac{10^{6}}{4\times 10^{5}}=2.5\] \[(1.05)^{x}=2.5\] Can you take the logs of both sides of the last equation?

OpenStudy (anonymous):

where do u get the 6 from for the power

OpenStudy (anonymous):

thats the million right

OpenStudy (anonymous):

man im so lost crap

OpenStudy (anonymous):

y is the x blank for the exponent

OpenStudy (kropot72):

X is the number of years from 1990 for the population to increase from 400,000 to 1,000,000.\[(1.05)^{x}=2.5\] Taking logs of both sides: \[x \times \ln 1.05=\ln 2.5\] \[x=\frac{\ln 2.5}{\ln 1.05}\]

OpenStudy (anonymous):

=( where did the 2.5 come from

OpenStudy (kropot72):

Check back on my previous posts. 2.5 is the ratio of 1 million to 400,000: \[\frac{10^{6}}{4\times 10^{5}}=2.5\]

OpenStudy (anonymous):

oh i see what u did u broke down the 400,000... oh ok

OpenStudy (anonymous):

oh ok then u divide that by the 1.05

OpenStudy (kropot72):

You divide the natural log of 2.5 by the natural log of 1.05 to find x:\[x=\frac{\ln 2.5}{\ln 1.05}\]

OpenStudy (anonymous):

2.38095

OpenStudy (anonymous):

the answer is "in between 2008 and 2009" btw

OpenStudy (kropot72):

No. You have just found 2.5/1.05. You need to use the 'ln' key on a calculator.

OpenStudy (anonymous):

.8675?

OpenStudy (anonymous):

im supposed to use excel for these problems not a calculator =(

OpenStudy (anonymous):

im seriously havin a hard time understanding this

OpenStudy (kropot72):

x = 18.78 years. If you add 18.78 to 1990 what year do you get?

OpenStudy (anonymous):

2008 i would so not be able to do a similar problem on my own

OpenStudy (anonymous):

so is there no easier way to do this?

OpenStudy (kropot72):

I just checked on my Excel. LN returns the natural logarithm.

OpenStudy (anonymous):

sorry just reviewing everything u posted

OpenStudy (kropot72):

You can also use ordinary logs (logs to base 10). The answer is exactly the same.

OpenStudy (anonymous):

ok thank u

OpenStudy (kropot72):

You're welcome :)